Two waves of wave leght m and M sounded together produce 7 beats per second what is the speed of the waves

Grade:11

1 Answers

Riddhish Bhalodia
8 years ago
Assuming their speed is same
The beats are caused by frequency difference
f_1 - f_2 = 7
and
f_1 = v/m \quad \textrm{and} \quad f_2 = v/M

hence by putting it in the equation we get
v = 7mM/(M-m)